A helicopter of mass 3500 kg is hovering. a) How much force must the rotors exert to do this? b) How much force in total must the rotors exert to make the helicopter accelerate upwards ate 1.5 m/s2 ? c) How much force in total must be exerted to allow the helicopter to accelerate downwards at -1.5 m/s2
1
Expert's answer
2019-12-02T09:52:40-0500



We write the equation of motion for each case


a)"F_1-m \cdot g=m \cdot a_1" where from

"F_1=m \cdot a_1+m \cdot g=m \cdot 0+9.81 \cdot 3500=34335[N]"


b) "F_2-m \cdot g=m \cdot a_2" where from

"F_2=m \cdot a_2+m \cdot g=3500 \cdot 1.5+9.81 \cdot 3500=39585[N]"

c)"F_3-m \cdot g=-m \cdot a_3" where from

"F_3=-m \cdot a_3+m \cdot g=-3500 \cdot 1.5+9.81 \cdot 3500=29085[N]"